### v4

# Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs

Abstract
Can LLMs pick up language structure from examples? Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures. However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages. In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall. We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaluate LLMs regardless of training data. We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks. These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones seen at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
Machine Learning, ICML
1 Introduction
Figure 1: We sample randomly generated languages to test LLMs by sampling deterministic finite automata (DFAs). (a) The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet. Accepted strings belong to the grammar, and rejected strings do not. We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the DFA-defined grammar, and (c) generate new strings following the grammar.
Contemporary LLMs have proven themselves to be highly sophisticated natural language completion models that demonstrate many properties of reasoning engines. This has prompted questions surrounding the true intelligence of these models, with some arguing that they possess inherent language learning capabilities (Millière, 2024). In this paper, we explore the question of whether LLMs have the reasoning capacity to understand the structure of a new language. Specifically, we are interested in problems where a model is given a small set of examples from a language and either generates a new sample or determines whether a new sample is from the language or not.
Some work suggests that LLMs broadly understand language structure because they are able to produce syntactically correct samples from languages they have not been trained on (Athiwaratkun et al., 2022), although LLMs’ performance on low-resource languages tends to be lower than their ability on higher resource languages (Bogin et al., 2023). On the other hand, some critics of LLMs argue that LLMs cannot possess an understanding of language structure as they have learned from data rather than possessing a priori universal grammar (Chomsky et al., 2023). We do not view learning from data as a fundamental limitation, but we are concerned with the possibility that a language model might only be able to understand linguistic structures similar to those it has seen in training data. To distinguish between these possibilities, we wish to evaluate LLMs on a benchmark that considers wholly novel languages, eliminating the possibility of dataset leakage.
Ideally, models would be tested on a set of language reasoning problems disjoint from data seen during training and validation. However, as training datasets for LLMs are generally closed and incredibly vast, human-generated problems in natural language are likely to at least partially overlap in syntax or concept with content LLMs have already seen, making probing a model for its ability to reason about the structure of natural languages nearly impossible. Additionally, the problem of determining whether two tasks are semantically identical is itself a nontrivial one. Therefore, ensuring that even an entirely novel invented problem is not a variation on a theme is intractable.
To circumvent this problem, we propose the following general approach: first we define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, we sample random languages from this space. By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with common ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
Experimental results using this approach suggest that contemporary LLMs possess less sophisticated language pattern recognition abilities than expected; underperforming basic, parameter-free n-gram language models on even the simplest languages. These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over language tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
In summary, we make the following contributions:
  1. 1.
    We introduce a benchmark for LLM language reasoning evaluation, disjoint from natural language web data.
  2. 2.
    We evaluate a suite of popular LLMs on instances of this benchmark and demonstrate that LLMs underperform compared to simple language model baselines.
  3. 3.
    We analyze the differences in behavior between these models, illustrating the influence of RLHF and chain-of-thought prompting on language reasoning capacity.
2 Related Work
2.1 Reasoning with LLMs
Reasoning is one of many “emergent abilities” (Wei et al., 2022a) possibly possessed by LLMs (Huang & Chang, 2022), although the nonlinear dependence of such emergent abilities on model size is disputed (Schaeffer et al., 2024). The chain-of-thought prompting technique (Wei et al., 2022b) has inspired a number of approaches to encourage the latent reasoning ability of models (Yao et al., 2023; Besta et al., 2024; Kojima et al., 2022), including neuro-symbolic methods (Hua & Zhang, 2022; Weir et al., 2023, 2024). Building on this, other work considers how to optimize exemplars used for in-context learning (Dong et al., 2022) and chain-of-thought prompting, known as “rationale refinement” (Liu et al., 2021; Fu et al., 2022). Problem-decomposition is also shown to be effective (Zhou et al., 2022; Khot et al., 2022).
2.2 LLM reasoning evaluation
LLM reasoning abilities are often tested on natural language benchmarks and commonly seen problems like arithmetic (Cobbe et al., 2021; Amini et al., 2019; Hendrycks et al., 2021), commonsense reasoning (Bhargava & Ng, 2022), and other, sometimes generative, tasks (Lake & Baroni, 2018; Pasupat & Liang, 2015; Lin et al., 2019) and task collections (Srivastava et al., 2022). LLMs have been shown to lack sufficient reasoning capability across a range of tasks including multi-step planning and complex inference (Valmeekam et al., 2022). Fan et al. (2023) introduce an LLM reasoning benchmark on algorithmic problems through NP-hard complexity, and Hazra et al. (2024) show that LLMs struggle to complete simple 3SAT problems. Patel et al. (2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, and Razeghi et al. (2022) similarly find that term frequency in training data impacts models’ in-context learning ability. In comparison to these, we explore the distinction described by Patel et al. (2021), but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs. This enables us to evaluate the ability of LLMs to reason about language.
2.3 Language Understanding and LLMs
LLMs can be quite adept at generating programs in general-purpose programming languages (Xu et al., 2022a). In contrast, adapting models to understand domain-specific languages (Mernik et al., 2005) introduces unique problems such as navigating idiosyncratic syntax and semantics, and leveraging sparse collections of sample language data. To address these challenges, researchers have considered how well general-purpose LLMs can use language reasoning skills to quickly understand rare or unseen DSLs with only a small set of exemplars (Joel et al., 2024). While most work in this vein focuses on semantic parsing for downstream applications (Lin et al., 2023), selecting exemplars (Zhao et al., 2021), and improving DSL recognition by leveraging more common languages (Bogin et al., 2023), experiments show strong baseline performance for LLM DSL recognition and parsing out-of-the-box (Wang et al., 2024), indicating that LLMs may possess emergent language reasoning abilities.
Related lines of work are compositional generalization (Xu et al., 2022b), which assesses models’ ability to organize known units into novel structures, and structural generalization (Yao & Koller, 2022), which assesses models’ ability to recognize new structures. Yao & Koller (2022) show that smaller language models like BART and T5 can struggle on these tasks, but to our knowledge there are not comprehensive experiments extending this line of work to LLMs.
2.4 Training transformers on Formal Languages
Vafa et al. (2025) frame world modeling as a latent DFA identification task, finding that transformers trained on DFA traces (of massive DFAs representing board games and city maps) do not reconstruct the underlying DFA. Akyürek et al. (2024) find that transformers trained on 4-12 state DFA transducer traces more effectively learn to in-context-learn regular languages than RNNs. Here, we evaluate much larger and more capable LLMs on much simpler 3-state DFAs. However, our goal is evaluating the abilities of foundation models (trained on a broad corpus rather than a tailored one), and is thus not strictly easier.
3 DFA Reasoning Tasks
3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA). Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Type 3 languages are also known as regular languages, which are recognized by regular expressions.
One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, i.e., the amount of working memory. [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious. 2-state DFAs have the property that their set of states is no larger than the output set $\{0,1\}$, and, therefore, do not have any hidden state. We thus explore 3-state DFAs, as this is the simplest nontrivial case.
Figure 2: An illustration of Chomsky’s hierarchy of languages, ranging from Type 0 to Type 3, which are defined by what formal models can recognize their grammars. In this work, we focus on the simplest language type in the hierarchy, regular grammars, which are recognized by deterministic finite automata (DFAs).
3.2 Sequence Completion Task
We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. In practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
To generate test cases for this task given a DFA, we first (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept. The task is to find a completion (not necessarily the same completion found in sampling) of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence. For details on sampling, see Appendix A.2.
We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random (each of which contains 30 example sequences and a distinct prefix), and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately, then computing a correctness metric as a fraction. We then average this metric over several sampled DFAs to produce our accuracy score.
3.3 Transducer Task
While the sequence completion task is the natural one that comes to mind as a basic language task, it has a difficulty-gap problem. Specifically, the issue is that many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model. Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.222 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token. As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model. Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences that end in a are in the language, making the problem trivial. To provide a more direct evaluation of non-world-modeling-based pattern recognition, we explore the Transducer task.
In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model. We call this a transducer task, as the DFA converts a sequence of inputs into a sequence of outputs. E.g., given the language does the string have an even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1. For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
This task is significantly more transparent than the sequence completion task as the model has access to intermediate outputs, an (imperfect) proxy for intermediate state. n
3.4 Baselines
l To contextualize LLM accuracies, waseline models with varying degrees of sophistication.
Sequence Completion Task
For the Sequence Completion task, we have three kinds of baseline.
  • RandomS baseline: produce a random string of length 5 characters. While this might seem redundant as it should have a success rate of 50%, in practice our rejection sampling approach (see Appendix A.2) leads to a slight bias towards DFAs with more accept states. This baseline measures that bias.
  • Common-SuffixS baseline: find the completion $s$ of length between 1 and 5 that maximizes ((((# of occurrences as a ${}\times|s|)$. This baseline does not take the distinct prefix into account, and instead tries to find a universal completion that will always end in an accept state for this language.
  • $n$-GramS baseline: we take the last $n-1$ characters of the distinct prefix and search to see if they appear in any of the example sequences at a position where the sequence following is an appropriate length to be a completion (at least 1 but at most 5). We then take a plurality vote among the completions and return this, breaking ties arbitrarily. If there are no matches, we return the result of $(n-1)$-GramS. Technically these cover more than $n$ characters, since the completion is often $>1$ character long; for simplicity, however, we keep the naming consistent with the Transducer baselines.
  • Brute-ForceS: take all possible DFAs with 3 states and 3 symbols. Filter for ones that accept all the example sequences. Then try all remaining DFAs on all $3^{5}$ possible 5-length completions and return the completion that the maximal number of DFAs accept, breaking ties arbitrarily.
Note that these baselines are entirely unparameterized and operate identically regardless of the underlying DFA. This makes them direct comparisons to using LLMs in in-context-learning. We do not consider BruteForceS to be a reasonable comparison due to its computational complexity, and instead consider it an upper bound on performance on this particular task.
Transducer Task
We have similar baselines for the Transducer task.
  • NullT baseline: for a given DFA, whichever of the following strategies produces a higher accuracy: always predict 0 or always predict 1.
  • $n$-GramT baseline: take the $n-1$ symbols ending at the end of the concatenated transducer sequence (e.g., for $n=5$ and the above example, this would be 1a0a). If that sequence does not appear elsewhere in the sequence, return the result of the $(n-1)$-GramT baseline. Otherwise, take the token that appears immediately after each occurrence. If there is a majority, return that, otherwise return the last example.
  • BruteForceT: take all possible DFAs with 3 states and 3 symbols. Filter them for ones that match the given transducer sequence. Take this set and predict the next token. Take a majority vote among these, returning 1 by default if there is no majority.
4 Experiments
We evaluated the open-source models Llama 3-8B, Llama 3-70B (AI@Meta, 2023), and Llama 3-8B-Instruct (AI@Meta, 2024), Mistral Nemo Minitron 8B (NVIDIA, 2024), Mistral Nemo Base 2407 (Mistral AI, 2024b) and Mistral Nemo Instruct 2407 (Mistral AI, 2024c), Gemma 7B (Google, 2024), and Falcon 7B (Almazrouei et al., 2023).
We also evaluated the open-source code models StarCoder2-15B (Lozhkov et al., 2024), Codestral-22B-v0.1 (Mistral AI, 2024a), Deepseek Coder 33B Instruct (Deepseek, 2024), Qwen2.5-Coder-7B, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), o1-preview (2024-09-12) (OpenAI et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024).
For each open source model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs. For GPT-4o and Claude, we evaluated on 30 DFAs due to computation costs. For o1-preview we evaluated on only 10 DFAs, and only on the Transducer task (which we felt was a better fit for a reasoning model). For gpt-3.5 and gpt-4o-mini, we evaluated on 100 DFAs. All models were evaluated with temperature 0, except o1-preview333o1-preview does not allow setting a non-default temperature..
For both tasks, we consider four prompting formats:
  • Basic provides no context, presenting the problem as a generic sequence generation or next-token prediction task, where output is provided immediately following the input, with no space to think.
  • More-Expl explains that the strings are generated from a DFA with 3 states, but is otherwise identical to Basic. This remains a sequence generation/next token prediction task.
  • COT provides the same information as More-Expl and additionally invokes chain-of-thought reasoning to help the model reason over the task. Here, the model is given space to reason before providing a tagged answer.
  • Red-Green casts the tasks as independent word problems that describe the underlying grammar structure without relying on world knowledge about DFAs and regular languages. It describes an N-state DFA as a house with N rooms, each of which has 3 portals that deterministically go to other rooms (or back to the same room), where the walls of each room are red or green (mirroring transducer output symbols 0 and 1). Similarly to COT, the model is given space to show work before providing a tagged answer.
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix F. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
Model Size IT? Code? Sequence Completion SR Transducer TR
Baselines
BruteForce 100.0 (99.9–100.0) 1 96.4 (96.2–96.7) 1
6-Gram 91.7 (91.0–92.4) 2 93.5 (93.1–93.9) 2
5-Gram 91.2 (90.4–91.9) 3 93.4 (93.0–93.7) 3
4-Gram 89.6 (88.7–90.4) 4 91.1 (90.6–91.6) 4
3-Gram 87.0 (86.1–87.8) 5 87.0 (86.4–87.6) 16
2-Gram 83.3 (82.2–84.2) 8 74.5 (73.6–75.3) 25
Common-Suffix 84.7 (83.6–85.6) 6
RandomS/NullT 53.3 (51.7–54.7) 26 68.9 (68.2–69.6) 26
Open Source Completion
llama3-8B 8.0B 73.8 (72.4–75.1) 18 87.5 (86.9–88.0) 14
llama3-70B 70.6B 71.4 (70.0–72.7) 23 87.7 (87.2–88.3) 12
llama3.1-8B-Instruct 8.0B 75.3 (74.0–76.6) 16 85.9 (85.3–86.5) 18
mistral-nemo-minitron-8B 8.4B 78.7 (77.5–79.8) 12 88.6 (88.0–89.1) 5
mistral-nemo-base-12B 12.2B 75.5 (74.3–76.6) 15 87.9 (87.4–88.4) 10
mistral-nemo-instruct-12B 12.2B 72.2 (70.9–73.4) 22 88.0 (87.5–88.5) 8
gemma-7b 8.5B 72.6 (71.3–73.7) 20 82.1 (81.4–82.7) 22
falcon-7b 7.2B 69.0 (67.6–70.2) 24 84.9 (84.3–85.5) 20
Open Source Code
starcoder2-15b 16.0B 73.5 (72.0–74.7) 19 87.7 (85.8–89.5) 13
codestral-22B 22.2B 78.0 (76.8–79.1) 13 86.6 (86.0–87.1) 17
deepseek-coder-33b-instruct 33.3B 76.7 (75.3–77.8) 14 85.6 (85.0–86.2) 19
qwen-2.5-coder-7B 7.6B 79.5 (78.4–80.5) 9 88.2 (87.6–88.7) 7
qwen-2.5-coder-instruct-7B 7.6B 79.5 (78.3–80.5) 10 88.3 (87.8–88.8) 6
qwen-2.5-coder-instruct-32B 32.8B 79.2 (78.0–80.3) 11 87.9 (87.4–88.4) 9
Proprietary
gpt-3.5-instruct ? 67.3 (63.1–71.5) 25 87.8 (85.9–89.6) 11
gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 27
gpt-4o-mini ? 72.4 (68.1–76.3) 21 79.8 (77.3–82.2) 23
gpt-4o ? 74.4 (69.9–78.6) 17 83.7 (80.1–86.9) 21
claude-3.5 ? 84.0 (79.3–88.4) 7 87.1 (83.9–90.2) 15
o1-preview ? 76.5 (69.4–84.3) 24
Table 1: Results for our experiments. We present model metadata alongside model results on both the Transducer and Sequence completion tasks. Each cell contains the mean performance across DFAs for the best-performing prompt (see Table 2 for details), with 95% confidence intervals of the mean in parentheses. “N/A” is used whenever the model returned an invalid result at least 25% of the time. (IT = Instruction-Tuned, TR/SR = Transducer/Sequence Completion rank, the ordinal rank of the given model on the given task.)
5 Results
Main results for all tasks are presented in Table 1. For all LLMs, we ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29. We then report the mean across DFAs and 95% bootstrap confidence intervals.
5.1 Sequence Completion
As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100accuracy in theory, as demonstrated by BruteForceS results. Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines. Still, we find that $n$-GramS heuristics tend to outperform LLMs.
As seen in Table 2, we find that giving the model the opportunity to logically reason about the prompt via chain-of-thought and present a conclusion has inconsistent results. Specifically, we find that BasicS is the best prompt for gpt-4o-mini, but not gpt-4o, where the best performing prompt is Red-GreenS. We find that claude-3.5 is entirely unable to follow the sequence completion prompts BasicS and More-ExplS, and performs best at the COTS prompt.
Additionally, we find that in this task, code-specific open-source models tend to perform better than sequence completion models, suggesting some generalized ability to produce strings from novel languages demonstrated by example. Overall, the relative performances of LLMs and prompts comport somewhat well to heuristics on which models and prompting strategies should work best. Nonetheless, LLMs underperform simple $n$-Gram heuristics.
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix E for details.
Model Basic More-Expl COT Red-Green
Sequence Completion
gpt-4o-mini 72.4 (68.1–76.3) 70.5 (66.4–74.6) 58.0 (53.4–62.4) 59.1 (54.9–63.2)
gpt-4o 72.1 (65.9–78.2) N/A 67.4 (60.8–73.8) 74.4 (69.9–78.6)
claude-3.5 N/A N/A 84.0 (79.3–88.4) 80.0 (74.9–85.2)
Transducer
gpt-4o-mini 79.8 (77.3–82.2) 76.7 (74.2–79.3) 65.2 (63.1–67.4) 74.5 (72.0–77.0)
gpt-4o 83.7 (80.1–86.9) 82.6 (79.1–85.9) 67.8 (63.1–72.3) 82.6 (78.8–86.3)
claude-3.5 86.9 (83.3–90.0) 87.1 (83.9–90.2) 76.4 (72.9–79.9) 82.9 (78.9–86.9)
Table 2: Results for models where we investigated multiple prompts (we only used Basic on other models). We bold the best prompt for each model. Non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
5.2 Transducer
Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ baseline only achieving 96.4% accuracy. Comparisons are still be valid as all models see the same fraction of unsolvable instances.
We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task. The relative performance of the models also does not correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o. Even within a model class we find no clear pattern: GPT-4o and o1-preview [^4] The particularly poor performance of o1-preview may be due to the model not supporting temperature 0. In other experiments, we found that GPT models with temperature 1 tended to perform poorly. See Appendix B for more details. are outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B. Coding models also demonstrate no advantage on this task.
The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models. To investigate that this is not specific to the Basic prompt, we investigate other prompts for chat models. As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task in all cases.
We conclude that LLMs are unable to perform well on the DFA transducer inference task. This failure cannot be attributed to a lack of world modeling, as $n$-GramT models do not construct world models. Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
5.3 Comparison of Benchmarks
Figure 3: Transducer and sequence completion results plotted against each other. Points represent the mean over several DFAs and intervals represent 95% confidence intervals. Points are colored by model type, with the best and worst model by each metric in each category labeled, as well as all baseline and proprietary models.
Figure 3 displays the relationship between model performance on the Sequence Completion and Transducer benchmarks. While at a high level, there is a positive correlation between the two, there are a few notable differences. For one, the Code models perform notably better than other open source models on Sequence Completion, but not on Transducer. Additionally, on Transducer, a ceiling on performance is observed, where LLMs cluster together between 3-GramT and 4-GramT performance; this clustering does not appear on the Sequence Completion benchmark.
5.4 Case Study: Sum Modulo 3 DFA
Correct Incorrect
Total 100% 100%
a is no-op 70% 73%
1b and 1c lead to 0 47% 57%
2-periodic 30% 47%
3-periodic 13% 13%
2 red rooms 7% 10%
Figure 4: Results on Sum Modulo 3 DFA. (a) MB=mistral-nemo-minitron-8B/BasicT, CR=claude-3.5/Red-GreenT. Venn diagram of errors (out of 1000). Labeled percentages are accuracies. (b) Results of qualitative analysis, out of 30 in both cases.
We investigate the transducer task on the DFA depicted in Figure 1. This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3. For this case study, we focus on the model/prompt combinations MB and CR, defined as
  • MB: mistral-nemo-minitron-8B/BasicT. Selected as it is the best performing combination overall.
  • CR: claude-3.5/Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
Figure 4a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA. Nearly all errors made by the 6-GramT model were also made by at least one LLM, while the two LLMs often made unique errors. While this task is better-known than most DFAs, all 3 models perform worse on this DFA than average.
We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer. Results of this analysis can be found in Figure 4b. We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output. Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state. These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed, as well as our finding that LLM performance decreases for tasks that are not solvable by $n$-Grams; see Appendix C for details.
[…]
Despite transformers’ high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
6 Conclusion
Our findings highlight significant weaknesses in large language models’ ability to generalize to entirely novel language reasoning problems, even simple ones solely involving next-token prediction on basic languages recognized by 3-state DFAs. These results, combined with that of previous work demonstrating that large language models can quite accurately perform a variety of language tasks, suggests that LLMs solve language problems via a mechanism distinct from general language reasoning ability. Our use of n-gram baselines and next-token prediction tasks allows us to exclude the possibility that the issue is primarily related to LLMs’ lack of world modeling or any inherent limitations of next-token prediction models. We believe our results suggest that LLMs have learned individual models of particular languages, but not a general theory of language.
Interestingly, in our transducer experiments, LLMs consistently perform better by directly predicting the next token than by explicitly reasoning through the problem. While our conclusions are limited by the finite nature of our prompt set, this suggests that they do, in fact, possess some latent understanding of language, but this understanding is inferior to basic n-gram models for $n>3$.
Many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages. More specifically, in tasks where there is a need to produce an output in a precise, atypical, format, we should be skeptical of the ability of LLMs to in-context-learn this format. For these tasks, it may be prudent to seek a new approach.
Impact Statement
Aside from the social consequences of this work as related to advancing the field of Machine Learning in general, this work has the goal of advancing the field of benchmarks in Machine Learning. While we view this as a positive objective, as it ensures that models are being evaluated fairly, it might have negative consequences insofar as benchmarking techniques might be best left unpublished to prevent deliberate or unintentional overfitting.
Appendix A Details on Sampling
A.1 Sampling of DFAs
We use rejection sampling to sample DFAs. Specifically, we uniformly sample a start state, then for each (source state, symbol) pair, we sample a post-transition state. We also randomly assign each state to be accept or reject with probability 50%. We then reject any DFA that has all accept or all reject states (so only DFAs with 1 or 2 accept states are allowed), or for which certain states are unreachable from the start state.
A.2 Sampling of Sequence Completion Tasks
To sample a sequence completion task, we first sample a DFA as described in Appendix A.1.
To sample a task instance, we sample example sequences and distinct prefix. Each example sequence is sampled uniformly from the space of $\{a,b,c\}^{10}$ and then rejected if the DFA does not accept the sequence. Our distinct prefix and completion are sampled uniformly from $\{a,b,c\}^{5}\times\{a,b,c\}^{5}$, and are rejected if the DFA does not accept the concatenation of the two, or if the prefix is the prefix of any of the previous sequences. We then discard the completion. If we, at any point, reject 50 sequences when attempting to sample a sequence or prefix, we return an error.
We run a “pilot” sampling for a DFA to ensure that it is valid, in which we sample an instance as described above. If there is an error in sampling this pilot instance, we reject the DFA. Otherwise, we proceed to sample our task instances. At this stage, if there is an error in sampling, we reject the instance rather than the DFA. This pilot sample rejection procedure leads to a slight bias towards 2-accept state DFAs over 1-accept state DFAs, as measured by the $\textsc{Random}_{S}$ baseline.
A.3 Sampling of Transducer Tasks
We sample a DFA as described in Appendix A.1, and then sample random sequences (30 in our experiments) and generate transducer traces. If every transducer trace ends with a 0 or every trace ends with a 1, we reject the DFA and resample.
Appendix B Results of o1-preview
We evaluated o1-preview on 10 DFAs, using 30 problem instances per DFA of the Transducer task, as in other Transducer experiments, and the $\textsc{Basic}_{T}$ prompt, as this is the most neutral prompt. Table 3 displays results on each DFA. Overall, while these results are not on a particularly large sample, they fairly definitively demonstrate that o1-preview does not achieve strikingly good performance on this task.
DFA o1-preview gpt-4o 6-Gram
1 25/30 27/30 26/30
2 23/29 24/30 25/30
3 19/30 23/30 28/30
4 22/30 23/30 28/30
5 29/29 30/30 30/30
6 19/30 24/30 30/30
7 17/29 23/30 25/30
8 23/30 25/30 26/30
9 21/30 28/30 30/30
10 29/30 29/30 30/30
Table 3: Results on each DFA. We find that in 7 cases, o1-preview underperforms gpt-4o, in 2 cases it gets the same number of instances wrong but provides a non-answer on an additional instance, and in 1 case it ties gpt-4o. In no cases does it outperform.
Appendix C Transducer results by difficulty class
Figure 5: Transducer results by difficulty class. We classify each DFA based on which of the baselines first achieves a score of 28/30 on the given instances. 6-Gram is excluded as it has very similar performance to 5-Gram. Each model’s best prompt results are plotted, with 95% confidence intervals, for all models with at least 100 DFAs; those with 10 or 30 had error bars too large to make this analysis useful.
Figure 5 displays results by difficulty level, as judged by the smallest $n$-Gram model that can solve a particular task. All models behave roughly monotonically, performing more poorly as difficulty increases. Additionally, we find that the best models continue to perform similarly to 4-Gram for tasks that 4-Gram does not perfectly solve.
Appendix D More details on Sum Modulo 3 DFA case study
Figure 6 depicts the results of the Sum Modulo 3 experiment, but filtered for two conditions. In the (a) condition, the trace ends in such a way that a 3-Gram model would be able to determine the output, and the (b) condition is the complement.
Figure 6: Results on Sum Modulo 3 DFA under trivial / nontrivial conditions. Percentages are accuracy numbers, and venn diagram is error counts. (a) In this condition, CR and the 6-GramT both get very high accuracies, with nearly all 6-GramT also being CR errors. MB does relatively poorly. (b) In this condition, models do significantly more poorly overall, with CR in particular performing worse than chance. Here, errors are more symmetric, with more 6-GramT errors that are not accounted for by either or both model, indicating that a larger fraction of both successes and failures in this condition are down to random chance.
Model BasicS Basic-CommasS
qwen-2.5-coder-7B 79.5 (78.4–80.5) 60.7 (59.3–62.1)
qwen-2.5-coder-instruct-7B 79.5 (78.3–80.5) 55.5 (54.0–56.9)
qwen-2.5-coder-instruct-32B 79.2 (78.0–80.3) 55.2 (53.7–56.7)
mistral-nemo-minitron-8B 78.7 (77.5–79.8) 59.3 (57.9–60.8)
codestral-22B 78.0 (76.8–79.1) 59.0 (57.5–60.3)
deepseek-coder-33b-instruct 76.7 (75.3–77.8) 54.9 (53.0–56.8)
mistral-nemo-base-12B 75.5 (74.3–76.6) 60.6 (59.1–62.2)
llama3.1-8B-Instruct 75.3 (74.0–76.6) 56.3 (54.4–58.1)
llama3-8B 73.8 (72.4–75.1) 61.5 (60.2–62.9)
starcoder2-15b 73.5 (72.0–74.7) 58.2 (56.7–59.8)
gemma-7b 72.6 (71.3–73.7) 54.0 (51.9–56.0)
gpt-4o-mini 72.4 (68.1–76.3) 64.1 (59.5–68.3)
mistral-nemo-instruct-12B 72.2 (70.9–73.4) 58.2 (56.4–59.8)
gpt-4o 72.1 (65.9–78.2) 66.8 (58.5–74.8)
llama3-70B 71.4 (70.0–72.7) 56.4 (54.7–58.0)
falcon-7b 69.0 (67.6–70.2) 56.1 (54.5–57.6)
gpt-3.5-instruct 67.3 (63.1–71.5) 52.3 (46.5–57.9)
claude-3.5 N/A N/A
gpt-3.5-chat N/A N/A
Table 4: Results on Sequence Completion Task. We compare BasicS to the comma-variant Basic-CommasS.
Appendix E Sequence Completion task prompt with Commas
To avoid tokenization differences with models, we also investigate a version of our Sequence Completion prompt that uses spaces and commas between the elements of the sequence. Unfortunately, results using this prompt were uniformly worse than results on the prompt without spaces and commas. Table 4 shows the results on a variety of models. All are worse with commas than without.
Appendix F Prompt Listings
Prompt $T$ $S$
Basic You are a sequence completion model. Output the next element of the sequence, and nothing else. <TRANSDUCER PREFIX>, The following strings come from an alien language that follows a simple grammar. Infer the alien grammar using the example strings. Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar. Output only the necessary suffix to complete the final string, and nothing else. <EXAMPLES> <PREFIX>
More-Expl You are a sequence completion model. The following sequence is generated from an unknown but consistent grammar. Identify the patterns within the sequence to determine its next element. Output the next element of the sequence, and nothing else. <TRANSDUCER PREFIX>, I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Output only the necessary suffix to complete the final string, and nothing else. <EXAMPLES> <PREFIX>
COT A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string. I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect the inputs and outputs into an input sequence and an output sequence. Infer the underlying DFA model to predict the next integer in the output sequence. Reason step by step, and then output the next output integer using <answer> tags, like <answer>0</answer>. Input sequence: <TRANSDUCER PREFIX> Output sequence: I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Reason step by step, and then output the next necessary suffix for this final string, <answer> tags, like <answer>ab</answer>. Given these valid strings: <EXAMPLES> Complete the following string: <PREFIX>
Red-Green ‘‘‘ You are in a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled A, B, and C. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way. In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *three* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. As you move through the house, at each time step you write down what portal you take and the color of the room you arrive (or stay) in. Based on your notes, predict what color room you will end up in after the last step. Tag your final answer like <answer>color</answer>. You walk through a portal labeled "<TRANSDUCER PREFIX>" and end up in a red room. ‘‘‘ You are outside a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled a, b, and c. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way. In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *3* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. Youve been into this house many times before. Each time, as you move through the house, you write down what series of portals you take and the color of the room you end up in. You have a collection of paths youve taken where youve ended up in a room with green walls, listed below. Given the final incomplete path at the bottom, write a series of up to 5 remaining steps that will cause you to end up in a room with green walls again. Tag your final answer like <answer>ab</answer>. Given these paths that end in a room with green walls: <EXAMPLES> Complete the following path: <PREFIX>
Table 5: Shortened summary of each prompt
F.1 Summaries
Table 5 contains summaries of each prompt.
F.2 Full example listings
F.2.1 BasicT
You are a sequence completion model. Output the next element of the sequence, and nothing else.
a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
F.2.2 More-ExplT
You are a sequence completion model. The following sequence is generated from an unknown but consistent grammar. Identify the patterns within the sequence to determine its next element. Output the next element of the sequence, and nothing else.
a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
F.2.3 COTT
A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string.
I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect the inputs and outputs into an input sequence and an output sequence. Infer the underlying DFA model to predict the next integer in the output sequence. Reason step by step, and then output the next output integer using <answer> tags, like <answer>0</answer>.
Input sequence: a, b, a, b, b, c, a, c, a, a, a, c, b, c, c, a, b, b, b, b, a, b, a, a, b, c, a, c, a, b
Output sequence: 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1,
F.2.4 Red-GreenT
You are in a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled A, B, and C. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way.
In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *three* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. As you move through the house, at each time step you write down what portal you take and the color of the room you arrive (or stay) in. Based on your notes, predict what color room you will end up in after the last step.
Tag your final answer like <answer>color</answer>.
You walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a red room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a red room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a red room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a ...
F.2.5 BasicS
The following strings come from an alien language that follows a simple grammar. Infer the alien grammar using the example strings. Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar. Output only the necessary suffix to complete the final string, and nothing else.
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
caacb
F.2.6 Basic-CommasS
The following strings come from an alien language that follows a simple grammar. Infer the alien grammar using the example strings. Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar. Output only the necessary suffix to complete the final string, and nothing else.
c, b, c, b, a, b, b, c, c, a
a, b, c, a, a, a, c, b, a, a
a, a, b, c, c, b, a, b, b, b
b, b, b, c, c, b, b, b, c, a
a, a, b, a, b, a, c, c, b, a
a, a, a, a, c, b, a, c, a, c
b, a, a, c, b, c, c, b, a, a
c, b, b, a, a, c, a, b, c, c
b, a, a, b, a, a, c, a, a, b
b, b, b, b, b, c, a, c, a, b
a, c, a, a, b, c, b, b, b, a
a, c, a, a, c, b, c, c, a, c
c, a, c, b, a, b, c, b, b, a
a, b, c, b, c, b, c, b, c, c
c, c, a, c, c, c, c, a, b, a
b, c, b, c, a, b, b, c, c, a
b, a, a, b, a, c, a, b, c, a
c, a, a, b, a, b, a, c, a, c
b, a, c, a, c, a, c, c, a, a
b, c, a, c, b, b, b, b, c, a
b, c, b, b, b, c, a, c, c, c
c, c, a, b, b, c, c, c, b, b
b, c, c, b, c, a, b, b, c, a
b, a, a, c, b, a, b, c, b, c
c, c, a, c, a, b, c, c, a, b
c, a, a, c, b, c, a, a, a, b
c, a, c, b, a, a, c, c, a, c
a, a, c, c, b, c, a, a, b, b
a, b, a, c, a, b, c, a, a, b
b, a, c, b, c, b, c, a, c, a
c, a, a, c, b,
F.2.7 More-ExplS
I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Output only the necessary suffix to complete the final string, and nothing else.
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
caacb
F.2.8 COTS
I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Reason step by step, and then output the next necessary suffix for this final string, <answer> tags, like <answer>ab</answer>.
Given these valid strings:
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
Complete the following string:
caacb
F.2.9 Red-GreenS
You are outside a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled a, b, and c. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way.
In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *3* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. Youve been into this house many times before. Each time, as you move through the house, you write down what series of portals you take and the color of the room you end up in. You have a collection of paths youve taken where youve ended up in a room with green walls, listed below. Given the final incomplete path at the bottom, write a series of up to 5 remaining steps that will cause you to end up in a room with green walls again.
Tag your final answer like <answer>ab</answer>.
Given these paths that end in a room with green walls:
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
Complete the following path:
caacb